OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (aakashsudhakar):

For all intents and purposes, since g(x) = -x + 3, it can be rewritten as g(x) = 3 - x. Therefore, we're given that f(x) = 9x - 2 and g(x) = 3 - x. f(g(x)) = f(3 - x) = 9(3 - x) - 2 = 27 - 9x - 2 = 25 - 9x f(g(x)) = 25 - 9x